OpenStudy (kendellfawn):

Mathematical Induction: 12 + 42 + 72 + ... + (3n - 2)2 = quantity n times quantity six n squared minus three n minus one all divided by two

OpenStudy (s4sensitiveandshy):

how would you start it ?? what is the first step ?

OpenStudy (s4sensitiveandshy):

\[12 + 42 + 72 + ... + (3n - 2)^2=\frac{ n(6n^2 -3n-1)}{ 2 }\] check if the statement is true for n=1 substitute all n's for 1 and then 2nd) step, assume that the statement is true for n=k (assumption) substitute n for k 3rd) step,based on the assumption prove that the statement is also true for n= k+1 substitute n for k+1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

have you ever done a proof by induction before?